# Can the 3 neutrino masses really be found using SN 1987A data?

### v1

@@ V.9 What if all three $m_{k}$ are all very small?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.4. showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$ In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst. However, once again we would not be able to identify the three lines about which the 30 points are clustering, as long as we are assuming very small neutrino masses. This means that the double bang scenario can be ruled out, given Fig.1.
Figure 4: If the 3 neutrino masses were very small as in the standard neutrino model, and none had $m_{j}^{2}<0,$ the three lines on $1/E^{2}$ versus t plot about which the 30 data points would cluster might look like Fig. 4, at least for those neutrinos emitted exactly at two times one second apart. But since neutrinos are in fact emitted at a continuum of times from $t=0$ to $t=1$ sec, there are many such line triplets will be the lines about which data points would cluster. Given 1 or 2 sec horizontal error bars, it clearly would be impossible to find any clustering of points about 3 specific lines, as is actually the case in Fig.1.
V.10 The preposterous size of the three $m^{2}_{j}$
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data, whether it be comological data or data from the KATRIN experiment implying $m_{\nu}<1eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the effective mass has $m_{k}^{2}<0$ the effective mass can be arbitrarily close to zero or even negative. In order to look for the separate $m_{k}^{2}$ in a direct mass experiment like KATRIN one needs to combine three spectra one for each mass with appopriate weights (Ehrlich2019a, ). Likewise, the constraint imposed on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided if one or more of the tachyon masses is negative gravitationaly. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, so that they have a repulsive gravitational effect. The best upper limit on the total sum of neutrino masses from cosmology is $\sum m_{\nu}<0.12eV/c^{2}$ using data from the Planck satellite and the Atacama Cosmology Telescope. Without allowing negative $m_{k}$ all three would need to obey $m_{k}<0.12eV/c^{2},$ a limit that does not apply if negative $m_{k}$ are allowed. In light of these considerations, the numerical values from the three mass fit in Fig.1 are not necessarily ruled out by either lab or cosmological data.
V.11 The tail of the neutrino emission spectrum
It was noted earlier that background data not associated with the SN 1987A burst would not be expected to lie on three straight lines. Fortunately, Bozza et al. has obtained four Kamiokande II data points having times $15>t>25$ seconds on the day of SN 1989A. According to the “tail” of the emissions spectrum in Fig. 3, such data would have a negligible chance of being part of the SN 1987A burst, and they are likely to be background. Background events would be significantly lower in energy than supernova events. Because of that energy difference those four data points for $15>t>25$ seconds in fact have only a $1\%$ chance of being consistent with that best fit in Fig. 1 according to chi square.
VI what does it all mean?
The eleven considerations of the previous section suggest that the three neutrino masses found in the three line fit (Fig. 1) are real, even if one of the masses has $m^{2}<0.$ This is not the first time the author has claimed that the three neutrino masses can be found from SN 1987A data. When the evidence was first published in 2013, the three neutrino masses were interpreted in terms of a $3+3$ model, i.e., three active-sterile mass doublets, where each doublet was split by one of three oscillation $\Delta m^{2}.$ (Ehrlich2013, ; Ehrlich2019, ) We now discuss a much more plausible explanation involving CPT violation.
VI.1 Possible Violation of CPT Symmetry
CPT symmetry is at the heart of the standard model and is well-tested by looking for measured particle-antiparticle mass differences. CPT violation upper limits have been found for a number of pairs of particles and their antiparticles including proton-antiproton, electron-positron and $K_{0}-\bar{K}_{0}.$ Thus, for example $|m(K_{0})-m(\bar{K}_{0})|<0.50\times 10^{-18}GeV,$ which can also be expressed in dimensionless terms:
$X_{CPT}=\frac{|m(K_{0})-m(\bar{K}_{0})|}{m(K_{0})}<10^{-18}$ (6)
However, if viewed as a constraint on the mass-squared separation, the bound appears rather weak for kaons, namely only $0.25eV^{2}.$ It is generally believed that neutrinos might offer the best opportunity of any particle for observing CPT violation if it exists. The usual way of looking for violations or setting upper limits on CPT violation for neutrinos has not been to find a neutrino antineutrino mass difference, but rather to compare oscillation parameters for neutrinos and antineutrinos. There are currently two accelerator-based long baseline experiments T2K-II, NOvA-II and a reactor based medium-baseline JUNO experiment capable of testing CPT symmetry in this manner. The synergy of these three experiments should be able to set an upper bound in the $\Delta m^{2}$ between neutrinos and antineutrinos of $5.3\times 10^{-5}eV^{2}$, extending the current bound twenty fold. (Ngoc2023, ) Furthermore, an analysis of data from short-baseline experiments, combining Gallium radioactive source (neutrino) data and reactor (antineutrino) data, pointed to a possible CPT-violating asymmetry in the effective mixing angles with a statistical significance of about $3.5\sigma.$
The strongest hint of evidence for CPT violation has been provided by Hisakazu Minakata who notes that accelerator neutrino experiments in the $\nu_{\mu}\rightarrow\nu_{e}$ channel and the $\nu_{\mu}\rightarrow\nu_{\mu}$ channel show the problem of “appearance-disappearance tension” at a high confidence level. (Minakata2025, ) Minakata suggests non-unitarity as a simple and natural way of resolving the tension. He has constructed a model with a unique solution that is consistent with data, and implies a large neutrino-antineutrino asymmetry (hence CPT violation). Minakata’s model bridges the two highest CL signatures of CPT violation, i.e. data from the BEST and LSND-MiniBooNE experiments, and the model implies a large neutrino-antineutrino asymmetry.
As the preceding duscussion suggests, neutrinos are the ideal place to look for CPT violation because
  • Strong hints of CPT violation have been seen
  • The extent of the violation would be larger than for any other particles
  • The mechanism behind the tiny neutrino masses is not fully explained by the Standard Model.
  • Unlike charged particles, the weak neutrino interactions mean experiments are less susceptible to ”fake” CPT-violating signals that can arise from electromagnetic interactions.
  • Neutrinos can travel vast distances through matter without interacting significantly. Experiments use this to their advantage, studying neutrinos over very long baselines and at high energies.
A CPT violation might involve a mass difference between neutrino and antineutrino, which would bring about $\nu-\bar{\nu}$ oscillations. A relevant parameter specifying the magnitude of any CPT violation here is $(m^{2}_{\nu}-m^{2}_{\bar{\nu}})/m^{2}_{\nu}=\Delta m^{2}/m^{2}.$ Since there are three neutrino mass states one would expect that the CPT symmetry breaking parameter $X_{CPT}=\Delta m^{2}/m^{2}$ would be the same for the three masses, i.e.,
$X_{CPT}=\Delta m_{1}^{2}/m_{1}^{2}=\Delta m_{2}^{2}/m_{2}^{2}=\Delta m_{3}^{2}/m_{3}^{2}$ (7)
[…]
$X_{CPT}=\Delta m_{atm}^{2}/m_{1}^{2}=\Delta m_{sol}^{2}/m_{2}^{2}=\Delta m_{sbl}^{2}/m_{3}^{2}$ (8)
where we have used for $m_{1},m_{2}$ and $m_{3}$ in Eq. 7 the three masses found in our three mass fit to the SN 1987A data. Thus, we have the following dimensionless result for the amount of CPT violation from Eq. 8: $X_{CPT}=5\times 10^{-6},$ for the three mass states. Note that this postulated value is an order of magnitude smaller than that attainable in the three experiments T2K-II, NOvA-II and JUNO (which explains why no CPT violation has yet been directly seen).
The possible CPT violation represented by Eq. 8 may follow from a suggestion by Eichorn and Shiffer’s involving quantum gravity. (Eichorn2025, ) In the early universe during the transplankian regime when $\nu-\bar{\nu}$ oscillations did take place and the neutrino masses and their separations became frozen in. There is another implication of CPT violation for neutrinos involving tachyons. Various theoretical models of quantum gravity suggest it could lead to both CPT violation and modifications to the speed of neutrinos, resulting in subluminal and superluminal propagation, and so one might expect one or more member of the mass triplet defined by Eq. 5 be tachyonic. i.e., have $m^{2}_{j}<0.$ The recognition that neutrinos might provide evidence for CPT violation goes back as far as Bruno Pontecorvo (Pontecorvo1957, ) in 1957 who first suggested $\nu\bar{\nu}$ oscillations which can only occur if they have unequal masses.
VI.2 Why is CPT violation more likely than $3+3$?
The author previously postulated a $3+3$ model (3 active-sterile neutrino pairs) to explain the three outlandish masses found in the fit to the SN 1987A data (Ehrlich2019, ). There are many advantages of the CPT violation explanation over the earlier $3+3$ model. Thus, we have seen neutrinos are a natural place for CPT violation to occur, strong hints it is occurring, and there are difficulties with sterile neutrinos. The primary difficulty with sterile neutrinos is the contradiction with results from numerous other highly sensitive experiments which have found no evidence for them, as well as tension with cosmological data. In addition, with three sterile neutrinos and oscillations allowed between sterile and active neutrinos in the $3+3$ model one has a $6\times 6$ PMNS mixing matrix which appears to be in conflict with the standard well-established $3\times 3$ matrix parameters that are in good agreement with data.
Figure 5: An artists work depicting a 3D reconstruction based on SN 1987A observations. Credit ESO/L Calcada
VI.3 Summary and a conclusive test
@@
This paper has made seven claims that some readers may consider very difficult to believe – see Table 1. Some of the seven claims will likely not be accepted as true without the strongest corroborating evidence.
@@ 2. No a priori upper limits exist on the three neutrino $m^{2}_{\nu,j}$
4. Negative search results cannot rule out tachyon’s existence
@@ 5. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
7. A likely explanation for these claims is CPT violation
Table 1: The seven claims made in this paper are based on the SN 1987A neutrino data. Regarding number 4, not only are negative search results incapable of proving tachyonic neutrinos non-existent, but they do not even necessarily make them less likely, but they may narrow the range of possible negative $m^{2}$ values.
We know SN 1987A was an asymmetric supernova (see Fig.5), with its expanding envelope and internal composition being lopsided rather than spherical. This spatial asymmetry was observed in the shape of the expanding envelope, and various other data. It is believed the asymmetry depicted likely arose from the progenitor star being a binary.
Figure 6: The asymmetric distribution of the debris of an explosion of the core of a 23-solar-mass progenitor star at 4.3 seconds after core bounce. The coloring codes mass density, blue being higher density. Credit: Argonne National Lab
Another particularly beautiful simulation of a supernova, this one lacking any symmetry whatsoever is shown in Fig. 6. However, it is not a mere spatial asymmetry that might corroborate the claims made in this paper, but rather a much deeper symmetry violation: the breaking of CPT symmetry and the evidence we have found supporting that violation by a dimensionless amount $X_{CPT}=5\times 10^{-5}.$ Usually, violations of the fundamental symmetries involve a decisive experiment. Thus, for parity (P) this was the C.S. Wu experiment, (Wu1957, ) for CP symmetry, this was the Fitch-Cronin experiment. (Fitch1964, ) Should a future experiment demonstrate a breaking of CPT symmetry for neutrinos by an amount $X_{CPT}=5\times 10^{-5}$ that would certainly be very strong corroboration of the seven claims. However, to find truly unambiguous evidence for all the claims, there is only one (and likely only one) decisive test, namely a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors.
The following concluding remarks represent one 88-year old’s supremely absurd optimism. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other six claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos, will definitely either confirm or reject the seven claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way to Earth from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say only 300 years). Were Betelgeuse to become a supernova, it would become incredibly bright, and even visible in daylight for many months. Go Betegeuse!
@@ Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments.
The author would be incredibly grateful to any readers
of this paper who find errors or poorly explained sections.

### v2

@@ IV.7 Suppose neutrinos all have $m<<1eV/c^{2}$?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.5. showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$
If the neutrino mass were in fact under the KATRIN current upper limit $(0.45eV/c^{)}$ the data points would all lie on lines that were virtually vertical and the data points would all lie in the shaded central area in Fig.1 having a 2 second time window. This is obviously far from reality, which suggests that their mass is not below the KATRIN limit. In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst, the preceding remains true, so once again we would not be able to identify the three lines about which the 30 points are clustering, unlike the actual data in Fig. 1. This means that the double bang scenario can be ruled out, given Fig.1.
Figure 5: The case of neutrinos having very small masses (which yield triplets of steeply sloped straight lines nearly overlapping) with a data point having a 2 sec wide horizontal error bar. Neutrinos are assumed emitted at $t=0$ and $t=1$ seconds. Given the size of the error bar, it would be impossible to find which of the six near vertical lines the data point is associated with (likewise for all the other line triplets for $0>t>1$ seconds – completely unlike what is actually the case in Fig.1.
IV.8 Consistency with direct mass experiments?
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data, whether it be cosmological data or data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit. KATRIN has further crucial relevance to that fit quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
IV.9 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum. (KATRIN2025, ). It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various masses and various weights. KATRIN then used a logarithmic grid of trial $m^{2}$ and $sin^{2}2\theta$ values that spans from $m^{2}=0.1eV^{2}$ to $1600eV^{2}$ and weight $sin^{2}2\theta$ from 0.001 to 0.5. At each grid point, the $\chi^{2}$ function is minimized with respect to three free parameters. One can then see how the goodness of fit varies over the grid, and how much of the grid is excluded based on a bad fit. By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid. This negative result was surprising because the Neutrino-4 group had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their neutrino oscillation experiment. Moreover, the KATRIN group is always very careful in their analysis of their data, so missing a signature of a sterile neutrino seems unlikely. Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
IV.10 Looking for the wrong kind of neutrinos?
A likely possibility is that KATRIN failed to see any $m>1eV/c^{2}$ neutrinos because they were looking for the wrong kind of neutrinos in their sterile neutrino search. They should have been looking for active $m>1eV/c^{2}$ neutrinos not sterile ones – based on what was found in our SN 1987A data analysis. Any $m>1eV/c^{2}$ active neutrinos unfortunately would not show up given the way the KATRIN sterile search was conducted. Thus, obviously, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. Rather they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three weights, $\alpha_{j},$ that is:
$R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})$ (6)
where $R_{\beta}$ is the usual beta decay formula for a given $m^{2}$, E is the electron energy and $E_{0}$ is the fitted endpoint energy for tritium beta decay. For the weights $\alpha_{j}$ only two are independent since $\sum\alpha_{j}=1,$ and the $\alpha_{3}$ for the tachyonic mass is obviously very small in view of the very large magnitude of $|m_{3}^{2}|.$
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the spectrum are found – doing this both for the three stipulated masses and away from them. Finally, for the functional form of the spectrum for the tachyonic third mass, it may be possible to use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. However, Ciborowski and Rembielinski have shown that for their tachyonic neutrino model a slightly more complicated formula is needed whose specific form depends on whether helicity or chirality coupling is used – see Eqs. 20-21 in ref. (Ciborowski1999, )
If KATRIN were to redo its analysis with these suggested changes it would be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$. A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest, while a negative result would doom the idea of obtaining neutrino masses from the SN 1987A data. The missing $2.70eV/c^{2}$ sterile neutrinos in KATRIN was the key clue here, in the proverbial “dog that didn’t bark.” category.
IV.11 Consistency with cosmological data?
An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data. That conflict on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided, if one of the tachyon masses is negative gravitationally. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect. The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses. (Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided if the mass sum were over flavor state masses not mass states.
In the recent literature the neutrino mass sum is considered to be the sum over mass state masses. However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states. Apart from historical usage there is a reason to view the sum as being over flavor state masses. Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old. At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor. For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each. Likewise for $e^{+}e^{-}$ collisions. In this case, the cosmological neutrino mass/energy density would be:
$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (7)
so, the appropriate constraint from Eq. 7 should be on the flavor state sum, and not the mass state sum, as usually assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
V Summary and the ultimate test
@@
This paper has made six claims that some readers may consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.9, and it finds evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. KATRIN likely can prove or reject all the above
Table 1: six claims made in this paper.
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors. These neutrinos will definitely either confirm or reject the seven claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, and easily visible in daylight. Go Betelgeuse!
@@ Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments. If KATRIN performs the decisive test suggested in section IV.9 while I am still on this planet, I thank Guido Drexlin and Diana Parno of the KATRIN Group.